# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v2

@@ Appendix B Proofs of RBK Convergence Theorems
Since the columns of $W_{2k}$ are orthonormal the random matrix $Z_{1}^{\top}W_{2k}$ can be redistributed as a single Gaussian random matrix $G_{2k}$ of size $k\times(2k)$ and with standard normal entries. By Lemma 3.16 of [37], we thus have for $k\geq 6$,
(59) $\mathbb{E}\left[\sigma_{\rm min}^{-2}(Z_{1}^{\top}L^{\top})\right]\leq\frac{20}{(\sqrt{2k}-\sqrt{k})^{2}\sigma_{2k}^{2}}\leq\frac{200}{k\sigma_{2k}^{2}}.$
Hence, by the independence between $(Z^{\top}L^{\top})^{+}$ and $Z^{\top}\tilde{r}$, we have
@@
$\mathbb{E}\,\|(Z^{\top}L^{\top})^{+}Z^{\top}\tilde{r}\|^{2}\leq\mathbb{E}\left[\sigma_{\min}^{-2}(Z_{1}^{\top}L^{\top})\right]\mathbb{E}\,\|Z_{2}^{\top}\tilde{r}\|^{2}\leq\frac{200}{\sigma_{2k}^{2}}\|\tilde{r}\|^{2}.$
@@
Noting that $\lim_{m\rightarrow\infty}\frac{\|r\|^{2}}{m}=\mathbb{E}\left[(a_{i}^{\top}x^{*}-b_{i})^{2}\right]=\|L^{\top}x^{*}-d\|^{2}=\|\tilde{r}\|^{2}$, we prove (20).
Proof B.5 (Proof of Corollary 3.3.).
@@
By Corollary 3.4 of [14], if $L$ has the polynomial spectral decay of order $\beta>1$, i.e., $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$, the dependence of $C_{n,k}$ on $n$ and $k$ in Theorem 3.2 can be eliminated when $k\leq n/2$. Furthermore, there is a constant $C=C(\beta)$ such that for any $k\leq n/2$, the linear convergence rate satisfies
$\lim_{m\rightarrow\infty}\alpha^{\rm RBK}\geq C\frac{k^{\beta}\sigma_{n}^{2}}{\|L\|_{F}^{2}}.$
@@
Regarding the convergence rate of mSGD, let us consider solving a fixed number of samples, say $m$, of $\{a_{i}^{\top},b_{i}\}$. It is demonstrated in [32] that, with a minibatch size of 1 and importance sampling, the corresponding convergence parameter can be at most $\alpha^{\rm SGD}\leq 1/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$ with $A_{m}=[a_{1},\dots,a_{m}]$ for convergent step sizes. When a larger minibatch size $k$ is used, [23] shows that the learning rate can be increased at most linearly, corresponding to a rate of at most $\alpha^{\rm mSGD}\leq k/\kappa_{\rm dem}(\frac{1}{m}A_{m}A_{m}^{\top})$. It follows by the random matrix theory [24, 43] that $\lim_{m\rightarrow\infty}\frac{1}{m}A_{m}A_{m}^{\top}=LL^{\top}$. Hence, when $m$ goes to infinity, the linear convergence rate of mSGD satisfies
$\lim_{m\rightarrow\infty}\alpha^{\rm mSGD}\leq\frac{k\sigma_{n}^{2}}{\|L\|_{F}^{2}}.$
@@ Appendix C Proofs of ReBlocK Convergence Theorems
(60) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$
@@
Since the ReBlocK iteration does not admit a simple orthogonal decomposition, we instead proceed by utilizing a bias-variance decomposition inspired by [12, 23, 18].
@@
(66) $\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]\leq(1-\alpha)\mathbb{E}\,\left\lVert d_{s}\right\rVert^{2}.$
@@
$\mathbb{E}\left[\left\lVert v_{s+1}\right\rVert^{2}|v_{s}\right]=$ $v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))^{2}\right]v_{t}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]$
$+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}$
$\leq$ $(1-\alpha)\left\lVert v_{s}\right\rVert^{2}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]$
$+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2},$
@@
Now, note that
(70) $A^{\top}W(S)r^{(\rho)}=A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r^{(\rho)}_{S}.$
Applying the singular value decomposition of $A_{S}$ and some elementary calculus, it is verified that regardless of the singular values of $A_{S}$,
(71) $\left\lVert A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}\right\rVert\leq\frac{1}{2\sqrt{\lambda k}}.$
The variance bound can thus be simplified as
(72) $\mathbb{E}\,\left\lVert v_{t}\right\rVert^{2}\leq\frac{1}{4\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$
Combining the bias bound (67) and the variance bound (72) using (63) yields the desired result.
@@
The result will follow from the appropriate application of Lemmas A.1, A.7, and C.1. To apply these lemmas it is first required to verify y $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)\preceq P_{RBK}(S)\preceq I$ for all $S$.
@@
(73) $x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert AA^{\top}+k\lambda I\right\rVert_{2}}>0,$
where the intermediate step uses the facts that $(A_{S}A_{S}^{\top}+k\lambda I)^{-1}\succeq(AA^{\top}+k\lambda I)^{-1}$, that row $i$ is chosen with probability $k/m$, and that $x^{\top}A_{S}^{\top}A_{S}x\geq x^{\top}a_{i}a_{i}^{\top}x$ when $i\in S$. It follows that $\overline{P}x\neq 0$ and so indeed $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$.
@@
The conditions of Lemma C.1 are also satisfied, the results of which provide the conditions for Lemma A.7 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $V=\frac{1}{2\lambda k\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$ Lemma A.7 then leads to the following bound for the tail averages:
(76) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x^{(\rho)}\right\rVert^{2}+\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}.$
Since $\rho=\mathbf{U}(m,k)$, the last term can be simplified using $\mathbb{E}\,_{S\sim\rho}\left\lVert r^{(\rho)}_{S}\right\rVert^{2}=k\left\lVert r^{(\rho)}\right\rVert^{2}/m$. The desired tail-averaged convergence bound follows.
The next step is to bound the condition number $\kappa(\overline{W})$, given that $\left\lVert a_{i}\right\rVert^{2}\leq N$ for all $i$. First note that the row norm bound implies
(77) $\left\lVert A_{S}A_{S}^{\top}\right\rVert=\left\lVert A_{S}^{\top}A_{S}\right\rVert=\left\lVert\sum_{i\in S}a_{i}a_{i}^{\top}\right\rVert\leq\sum_{i\in S}\left\lVert a_{i}a_{i}^{\top}\right\rVert\leq Nk.$
As a result,
(78) $\lambda kI\preceq A_{S}A_{S}^{\top}+\lambda kI\preceq(N+\lambda)kI.$
@@
(79) $\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\succeq\frac{1}{(N+\lambda)m}I,$
@@
(81) $\kappa(\overline{W})\leq 1+\frac{N}{\lambda},$
and the final claim of the theorem follows by Lemma A.9.
@@ Appendix D Noisy Linear Least-squares
In this section we consider the special case in which the inconsistency in the problem is generated by zero-mean noise. This case is more general than the case of Gaussian data, but less general than the case of arbitrary inconsistency. We define such a problem to have each row independently generated as
(82) $a_{i}\sim\mathbf{A},\;z_{i}\sim\mathbf{Z},\;b_{i}=a_{i}^{\top}y+z_{i},$
where $\mathbf{A}$ is arbitrary, $\mathbf{Z}$ satisfies $\mathbb{E}_{z\sim\mathbf{Z}}\left[z\right]=0$, and $y\in\mathbb{R}^{n}$. This scenario might arise, for example, if the entries of $b$ are calculated via noisy measurements of $y$ along the directions $a_{i}$.
For such problems with increasing numbers of rows, both RBK and ReBlocK converge (in a Monte Carlo sense) to the ordinary least-squares solution rather than a weighted solution. This can be viewed as an advancement over previous results on noisy linear systems, such as [30], which show only convergence to a finite horizon even for arbitrarily large noisy systems. The advance comes from treating the noise model explicitly and applying tail averaging to the algorithm.
@@
Consider the RBK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S})^{+}$ and $\rho=\mathbf{U}(m,k)$, applied to the noisy linear least-squares problem (82). Then the results of Theorem 3.1 hold and
(83) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
@@
Consider the ReBlocK-U algorithm, namely Algorithm 1 with $M(A_{S})=(A_{S}A_{S}^{\top}+\lambda kI)^{-1}$ and $\rho=\mathbf{U}(m,k)$, applied to the noisy linear least-squares problem (82). Then the results of Theorem 4.1 hold and
(84) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$
@@
Our strategy will be to first show that $x^{*}$ converges to $y$, then show that $x^{(\rho)}$ converges to $x^{*}$.
For the first part, note that for fixed $m$, the dataset $A,b$ can be viewed as a finite sample of the underlying statistical linear regression problem
(85) $y=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{a\sim\mathbf{A},z\sim\mathbf{Z}}\left[(a^{\top}(x-y)-z)^{2}\right].$
Furthermore, the ordinary least-squares solution $x^{*}$ is the corresponding empirical risk minimizer. It is a standard result in asymptotic statistics that as the number of samples $m$ grows, the empirical minimizer $x^{*}$ converges to the true solution $y$ [25, 44].
For the second part, recall that the weighted solution $x^{(\rho)}$ is characterized by
(86) $\overline{P}x^{(\rho)}=A^{\top}\overline{W}b=\overline{P}x^{*}+A^{\top}\overline{W}r.$
It has been verified in the proofs of Theorems 3.1 and 4.1 that $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. Combining this with the results of Lemma A.3 implies that $x^{*},x^{(\rho)}\in\operatorname{range}(\overline{P})$. Applying the pseudoinverse of $\overline{P}^{+}$ on the left thus yields
(87) $x^{(\rho)}=x^{*}+\overline{P}^{+}A^{\top}\overline{W}r,$
and it then suffices to show that
(88) $\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r=0.$
As $m\rightarrow\infty$, the distribution of $k$ rows sampled uniformly at random from $A$ converges to the distribution of $k$ rows sampled independently from $\mathbf{A}$. Similarly, the distribution of $k$ entries of the residual $r=Ax^{*}-b$ sampled uniformly from $A,b$ converges to the distribution of $k$ values sampled independently from $\mathbf{Z}$. By the noisy construction, these residual values are independent from each other as well as from the corresponding rows of $A$. Call the resulting distributions $\mathbf{A}^{k}$ and $\mathbf{Z}^{k}$. It then holds
(89) $\lim_{m\rightarrow\infty}\overline{P}=\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[P(A_{S})\right]=\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[P(A_{S})\right]\coloneqq\hat{P},$
$\lim_{m\rightarrow\infty}\overline{P}^{+}A^{\top}\overline{W}r$ $=\hat{P}^{+}\lim_{m\rightarrow\infty}\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[A_{S}^{\top}M(A_{S})r_{S}\right]$
$=\hat{P}^{+}\mathbb{E}_{A_{S}\sim\mathbf{A}^{k}}\left[A_{S}^{\top}M(A_{S})\right]\mathbb{E}_{r_{S}\sim\mathbf{Z}^{k}}\left[r_{S}\right]=0.$
It follows that
(90) $\lim_{m\rightarrow\infty}x^{(\rho)}=\lim_{m\rightarrow\infty}x^{*}=y.$

### v3

@@ Appendix B Proofs of RBK Convergence Theorems
Since the columns of $W_{2k}$ are orthonormal the random matrix $Z_{t}^{1}W_{2k}$ can be redistributed as a single Gaussian random matrix $G_{2k}$ of size $k\times(2k)$ and with standard normal entries. By Lemma 3.16 of [rebrova2021block], we thus have for $k\geq 6$,
(64) $\mathbb{E}\left[\sigma_{\rm min}^{-2}(Z_{t}^{1}L_{n}^{\top})\right]\leq\frac{20}{(\sqrt{2k}-\sqrt{k})^{2}\sigma_{2k}^{2}}\leq\frac{200}{k\sigma_{2k}^{2}}.$
Hence, by the independence between $Z_{t}^{1}$ and $Z_{t}^{2}$ we have
@@
(65) $\mathbb{E}\,\left\lVert(Z_{t}^{\top}L_{a}^{\top})^{+}Z_{t}^{\top}\tilde{r}\right\rVert^{2}=\mathbb{E}\left[\sigma_{\rm min}^{-2}(Z_{t}^{1}L_{n}^{\top})\right]\mathbb{E}\,\left\lVert Z_{t}^{2}\tilde{r}_{n+1}\right\rVert^{2}\leq\frac{200}{\sigma_{2k}^{2}}\cdot\|\tilde{r}\|^{2}.$
@@
Noting that $\mathbb{E}_{[a^{\top}\;b]\sim\mathcal{N}(0,Q)}\left[(a^{\top}x^{*}-b)^{2}\right]=\|L_{a}^{\top}x^{*}-\ell_{b}\|^{2}=\|\tilde{r}\|^{2}$, we prove (21).
Proof B.5 (Proof of Corollary 4.2).
@@
By Corollary 3.4 of [derezinski2024sharp], if $L_{n}$ has the polynomial spectral decay of order $\beta>1$, i.e., $\sigma_{i}^{2}\leq i^{-\beta}\sigma_{1}^{2}$ for all $i$, the dependence of $C_{n,k}$ on $n$ and $k$ in Theorem 4.1 can be eliminated when $k\leq n/2$. Furthermore, there is a constant $C=C(\beta)$ such that for any $k\leq n/2$, the linear convergence rate satisfies
$\alpha^{\rm RBK}\geq C\frac{k^{\beta}\sigma_{n}^{2}}{\|L_{n}\|_{F}^{2}}.$
@@
Regarding the convergence rate of mSGD, first consider a finite matrix $A$. It is demonstrated in [needell2014stochastic] that, with a minibatch size of 1 and importance sampling, the corresponding convergence parameter can be at most $\alpha^{\rm SGD}\leq\kappa_{\rm dem}^{-2}(A)$ for convergent step sizes. It is straightforward to generalize this result to the statistical problem Eq. 19, where the corresponding convergence bound is $\alpha^{\rm SGD}\leq\kappa_{\rm dem}^{-2}(L_{n})$. When a larger minibatch size $k$ is used, [jain2018parallelizing] shows that the learning rate can be increased at most linearly, corresponding to a rate of at most
$\alpha^{\rm mSGD}\leq k\kappa_{\rm dem}^{-2}(L_{n})=\frac{k\sigma_{n}^{2}}{\|L_{n}\|_{F}^{2}}.$
@@ Appendix C Proofs of ReBlocK Convergence Theorems
(66) $\mathbb{E}\,\left\lVert x_{t}-x^{(\rho)}\right\rVert^{2}\leq 2(1-\alpha)^{t}\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}+\frac{2}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r^{(\rho)}_{S}\right\rVert^{2}.$
@@
Since the ReBlocK iteration does not admit a simple orthogonal decomposition, we instead proceed by utilizing a bias-variance decomposition inspired by [defossez2015averaged, jain2018parallelizing, epperly2024randomizedkaczmarztailaveraging].
@@
(72) $\mathbb{E}\left[\left\lVert d_{s+1}^{2}\right\rVert|d_{s}\right]\leq(1-\alpha)\left\lVert d_{s}\right\rVert^{2}.$
@@
$\mathbb{E}\left[\left\lVert v_{s+1}\right\rVert^{2}|v_{s}\right]$ $=v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))^{2}\right]v_{t}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2}$
$\leq(1-\alpha)\left\lVert v_{s}\right\rVert^{2}+2v_{s}^{\top}\mathbb{E}_{S\sim\rho}\left[(I-P(S))A^{\top}W(S)r^{(\rho)}\right]+\mathbb{E}\,_{S\sim\rho}\left\lVert A^{\top}W(S)r^{(\rho)}\right\rVert^{2},$
@@
Noting that
(76) $A^{\top}W(S)r^{(\rho)}=A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r^{(\rho)}_{S}$,
we can combine the bias bound (73) and the variance bound (75) using (69) to obtain the desired result.
@@
The result will follow from the appropriate application of Lemmas A.1, A.7, and C.1. To apply these lemmas it is first required to verify $\overline{P}\preceq I$ and $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$. For the first result, it suffices to observe that $P(S)\preceq P_{RBK}(S)\preceq I$ for all $S$.
@@
(77) $x^{\top}\overline{P}x=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[x^{\top}A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}A_{S}x\right]\geq\frac{k}{m}\cdot\frac{(a_{i}^{\top}x)^{2}}{\left\lVert A^{\top}A+\lambda kI\right\rVert_{2}}>0,$
where the intermediate step uses the facts that $\left\lVert A_{S}A_{S}^{\top}+\lambda kI\right\rVert=\left\lVert A_{S}^{\top}A_{S}+\lambda kI\right\rVert\leq\left\lVert A^{\top}A+\lambda kI\right\rVert$, that row $i$ is chosen with probability $k/m$, and that $x^{\top}A_{S}^{\top}A_{S}x\geq x^{\top}a_{i}a_{i}^{\top}x$ when $i\in S$. It follows that $\overline{P}x\neq 0$ and so indeed $\operatorname{range}(\overline{P})=\operatorname{range}(A^{\top})$.
@@
The conditions of Lemma C.1 are also satisfied, the results of which provide the conditions for Lemma A.7 with $B=\left\lVert x_{0}-x^{(\rho)}\right\rVert^{2}$, $\tilde{V}=\frac{2}{\alpha}\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r^{(\rho)}_{S}\right\rVert^{2}.$ Lemma A.7 then yields the desired bound for the tail averages.
The next step is to bound the condition number $\kappa(\overline{W})$. Recall from the proof of Theorem 3.1 that $\left\lVert A_{S}A_{S}^{\top}\right\rVert\leq k\cdot\max_{i}\left\lVert a_{i}\right\rVert^{2}$, so that
(80) $\lambda kI\preceq A_{S}A_{S}^{\top}+\lambda kI\preceq(\max_{i}\left\lVert a_{i}\right\rVert^{2}+\lambda)\cdot kI.$
@@
(81) $\overline{W}=\mathbb{E}_{S\sim\mathbf{U}(m,k)}\left[I_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}I_{S}\right]\succeq\frac{1}{(\max_{i}\left\lVert a_{i}\right\rVert^{2}+\lambda)m}I,$
@@
(83) $\kappa(\overline{W})\leq 1+\frac{1}{\lambda}\cdot\max_{i}\left\lVert a_{i}\right\rVert^{2}.$
The bounds on the weighted residual and the bias follow from Lemma A.9 with the bias bound taking the standard form
(84) $\|x^{(\rho)}-x^{*}\|\leq\sqrt{\kappa(\overline{W})-1}\cdot\|A^{+}\|\cdot\|r\|.$
Since $\max_{i}\left\lVert a_{i}\right\rVert^{2}\leq\left\lVert A\right\rVert^{2}$, the combination of Eqs. 84 and 83 implies the looser bound
(85) $\|x^{(\rho)}-x^{*}\|\leq\frac{1}{\sqrt{\lambda}}\cdot\kappa(A)\cdot\left\lVert r\right\rVert.$
Lastly, consider the variance $V=\mathbb{E}\|A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r_{S}^{(\rho)}\|^{2}$. Since for $\sigma\geq 0$ one has $\sigma/(\sigma+\lambda k)^{2}\leq 1/(4\lambda k)$, it holds
$\big\|A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}\big\|^{2}=\lambda_{\max}\!\Big((A_{S}A_{S}^{\top})(A_{S}A_{S}^{\top}+\lambda kI)^{-2}\Big)\ \leq\ \frac{1}{4\lambda k}.$
Thus
[…]
@@ Appendix D Noisy Linear Least-squares
In this section we consider the special case in which the inconsistency in the problem is generated by zero-mean noise. This case is more general than the case of Gaussian data, but less general than the case of arbitrary inconsistency. In particular, we consider the statistical linear regression problem
(86) $x^{*}=\underset{x\in\mathbb{R}^{n}}{\operatorname{argmin}\;}\mathbb{E}_{a\sim\mathbf{A},z\sim\mathbf{Z}}\left[(a^{\top}x-(a^{\top}x^{*}-z))^{2}\right]$
where $\mathbf{A}$ is arbitrary, $\mathbf{Z}$ satisfies $\mathbb{E}_{z\sim\mathbf{Z}}\left[z\right]=0$, and $x^{*}\in\mathbb{R}^{n}$. This scenario might arise, for example, when making noisy measurements of $x^{*}$ along the directions $a$.
The RBK-U and ReBlocK-U algorithms are applied to Eq. 86 by generating $k$ independent samples per iteration from the data distributions $\mathbf{A},\mathbf{Z}$. The result is that RBK and ReBlocK both converge (in a Monte Carlo sense) to the ordinary least-squares solution rather than a weighted solution. This can be viewed as an advancement over previous results on noisy linear systems, such as [needell2010randomized], which show only convergence to a finite horizon even for arbitrarily large noisy systems. The advance comes from treating the noise model explicitly and applying tail averaging to the algorithm.
@@
Consider the RBK-U algorithm applied to the noisy linear least-squares problem (86). Then the results of Theorem 3.1 hold and $x^{(\rho)}=x^{*}$.
@@
Consider the ReBlocK-U algorithm applied to the noisy linear least-squares problem (86). Then the results of Theorem 5.1 hold and $x^{(\rho)}=x^{*}$.
@@
Recall that the weighted solution $x^{(\rho)}$ is characterized by
(87) $\overline{P}x^{(\rho)}=A^{\top}\overline{W}b=\overline{P}x^{*}+A^{\top}\overline{W}r=\overline{P}x^{*}+\mathbb{E}\left[A_{S}^{\top}M(S)r_{S}\right]$
where $r_{S}=A_{S}x^{*}-b_{S}$. Based on the definition of the noisy equation Eq. 86, each entry of the residual satisfies $r_{i}=a_{i}^{\top}x^{*}-(a_{i}^{\top}x^{*}-z_{i})=z_{i}$ and is thus distributed exactly according to $\mathbf{Z}$. In particular each entry of $r_{S}$ is independent from $A_{S}$ and the other entries of $r_{S}$ and has a mean of zero. It follows that
(88) $\mathbb{E}\left[A_{S}^{\top}M(S)r_{S}\right]=\mathbb{E}\left[A_{S}^{\top}M(S)\right]\mathbb{E}\left[r_{S}\right]=0,$
which completes the proof that $x^{(\rho)}=x^{*}$.